OpenStudy (anonymous):

Consider the expansion was (m+n)^25. What is the exponent of n in the 26th form

OpenStudy (amistre64):

proper use of the binomial thrm will aid us in this solution.

OpenStudy (amistre64):

recall for simplicity (a+b)^0 = a^0 + b^0 (a+b)^1 = a^1 + b^1 (a+b)^2 = a^2 +2ab+ b^2 (a+b)^3 = a^3 +3a^2b +3ab^2+ b^3 see if you can detemrine a pattern